OpenStudy (anonymous):

Suppose the diameter of a circle is 8. What is its area?

OpenStudy (study_buddy99):

I'm going to recommend you listen to this song: In the meantime, \[\pi*r^2\]

OpenStudy (study_buddy99):

I'll post the song in a second...

OpenStudy (study_buddy99):

so back to the problem: diameter is all the way across and radius is half way so pi=3.14 and your radius is 4 (half of 8) and 4^2= 16 because 4*4=16 so finally take 3.14*16=50.24 (your answer) any questions?

OpenStudy (study_buddy99):

https://www.youtube.com/watch?v=-z4SUypJZxo this song willl help SOO SOO much